# QuEST: Low-bit Diffusion Model Quantization via Efficient Selective Finetuning

### v2

@@ Abstract
Diffusion models have achieved remarkable success in image generation tasks, yet their practical deployment is restrained by the high memory and time consumption. While quantization paves a way for diffusion model compression and acceleration, existing methods totally fail when the models are quantized to low-bits. In this paper, we unravel three properties in quantized diffusion models that compromise the efficacy of current methods: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules. To alleviate the intensified low-bit quantization difficulty stemming from the distribution imbalance, we propose finetuning the quantized model to better adapt to the activation distribution. Building on this idea, we identify two critical types of quantized layers: those holding vital temporal information and those sensitive to reduced bit-width, and finetune them to mitigate performance degradation with efficiency. We empirically verify that our approach modifies the activation distribution and provides meaningful temporal information, facilitating easier and more accurate quantization. Our method is evaluated over three high-resolution image generation tasks and achieves state-of-the-art performance under various bit-width settings, as well as being the first method to generate readable images on full 4-bit (i.e. W4A4) Stable Diffusion. Code is been made publicly available at URL .
Machine Learning, ICML
@@ 1 Introduction
Diffusion models (Ho et al., 2020; Dhariwal & Nichol, 2021; Rombach et al., 2021) have achieved great success in the field of image generation recently. Given the high capacity and flexibility of diffusion models, there still remains two challenges that limit their efficiency (Croitoru et al., 2023). The first obstacle is the time-consuming denoising process which requires hundreds or thousands of time steps that slows down the inference speed drastically. The other one is the increasing model size driven by the need of better image fidelity and higher image resolutions. Both factors contribute to considerable latency and increased computational requirements, impeding the application of diffusion models to real-world settings where both time and computational power are restricted.
@@
Neural network quantization offers a feasible solution for accelerating inference speed and reducing memory consumption simultaneously (Gholami et al., 2022). It aims to compress high-bit model parameters into low-bit approximations with negligible performance degradation. For example, using 4-bit weight and 4-bit activation quantization can achieve up to 8$\times$ inference time speedup and 8$\times$ memory reduction theoretically (Liang et al., 2021). The benefits will increase consistently when the bit-widths are further lowered. Hence, low-bit quantization of diffusion models emerges as a viable approach for enhancing their efficiency.
@@
However, existing diffusion model quantization approaches fail under low-bit settings. These methods focus either on timestep-aware calibration data construction (Shang et al., 2023; Li et al., 2023a) or quantization noise correction (He et al., 2023b), targeting at tailoring existing quantization techniques to the properties of diffusion models that are distinct from other model types such as CNNs (Pilipović et al., 2018) and ViTs (Chitty-Venkata et al., 2023). Such approaches overlook the intrinsic diffusion model mechanisms that are related with quantization, leading to inconsistencies between method deployment and model characteristics. Therefore, we intend to identify unique features in diffusion models that are associated with quantization.
@@
In this paper, we reveal three properties of quantized diffusion models that hinder effective quantization: The activation distributions are likely to be imbalanced, where the majority of values are close to 0 but the other values are large and appear inconsistently. Conventional quantization methods tend to either approximate either the large and sparse values, inadequately estimating the numerous small values, or focus on small values while overlooking the large ones, thereby impeding the reduction of quantization error. Diffusion models exhibit varying functions at distinct time steps (Choi et al., 2022), such as forming object outlines at early time steps and constructing details latter. Thus, preserving accurate temporal information is important during quantization. However, current methods struggle to represent extensive time steps with limited bit-width under low-bit quantization, leading to similar time embeddings across different time steps and results in error accumulation. Diffusion models typically have complex network architectures, incorporating convolutional layers and linear layers that construct various kinds of attention blocks (Dosovitskiy et al., 2021) and residual blocks (He et al., 2015). Whereas previous works consider each model component as equally important and apply quantization uniformly, we reveal that certain layers are particularly sensitive to perturbations from quantization, while others are more resilient.
@@
Addressing the revealed properties, we propose a novel quantization approach for diffusion models, termed QuEST (Quantization via Efficient Selective FineTuning). Confronting property , we first theoretically justify that weight finetuning can enhance model robustness towards large perturbations under low-bit settings for reducing quantization error, whereas previous methods struggle to find appropriate quantization parameters. Then we empirically show that by finetuning the model, the activation distributions are actually modified to be more amenable to quantization. Following the idea of weight finetuning, we identify two types of layers that are responsible for performance degradation: Time embedding layers exhibiting property and attention-related layers corresponding to property . Thus, we selectively and progressively finetune these small layer subsets in conjunction with all activation quantization parameters, as illustrated in Figure 1(b). With less than 10% of the total parameters involved, QuEST not only substantially enhances low-bit quantized model performance, but is also notably time-efficient and requires only supervision from the full-precision model in a data-free manner. To preserve the essential time information, we further introduce an efficient Time-aware Activation Quantizer (shown in Figure 1(a)) that can flexibly adapt to the varying activation distributions over the numerous time steps. Our contributions are summarized as follows:
@@
  • We uncover and validate three properties that present challenges to low-bit diffusion model quantization. We also theoretically discuss why previous methods are expected to fail and justify the sufficiency of weight finetuning under low-bit settings.
@@
  • Addressing the identified properties, we introduce QuEST, an innovative finetuning strategy that trains the model selectively and progressively, achieving low-bit quantization capability with high time and memory efficiency. An efficient activation quantizer is also proposed for effective time information estimation. Practically, our method can be executed in a data-free and parameter-efficient manner.
@@
Figure 1: Framework of the proposed method at time step $t$. $s_{l}^{t}$ represents the quantization parameters of layer $l$ at time $t$, $Q(\textbf{w}_{l})$ represents the quantized weight of layer $l$. (a) In a quantized layer, the scaling factor $s_{l}^{t}$ is updated while the others are frozen at time $t$. (b) $e(t)$ and $x_{t}$ are the input time embedding and input image representation respectively. The model layers’ weights are optimized selectively and the quantization parameters are all updated. While different selected layers are learned progressively, the corresponding weight and scaling factors are updated simultaneously using intermediate activation alignment and task loss guidance.
2 Related Work
2.1 Diffusion Model Inference
Diffusion models (Ho et al., 2020; Rombach et al., 2021) generate samples via a learned iterative denoising process. During inference, the initial input is sampled from a Gaussian distribution: $x_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and the final output $x_{0}$ is obtained by removing noises from $x_{T}$ subsequently using:
In this Markov Chain, $\tilde{\boldsymbol{\mu}}_{\theta,t}$ and $\tilde{\beta_{t}}$ are calculated from the model’s output. For a typical diffusion model, the denoising stage requires hundreds to thousands of iterations. This laborious sequential process makes efficient inference and high-resolution generation extremely challenging. Latent Diffusion Models (LDMs) (Rombach et al., 2021) were developed to overcome the challenges. They encode the data into a latent space and then apply the diffusion process within this space, enabling high-resolution image generation with fewer time steps. The output images are produced by decoding the latent representations. LDM’s concept and network quantization complement each other, mutually enhancing the efficiency of diffusion models.
Practically, diffusion models typically adopt a UNet architecture (Ronneberger et al., 2015), with some modules substituted by attention blocks. In a latent diffusion model, an encoder and a decoder are incorporated for projecting images into latent representations and subsequently generating images from the latent space, respectively. Usually encoders and decoders are lightweight and computationally inexpensive, so our focus is on quantizing the UNets in LDMs, in alignment with other works (as described below).
Table 1: Comparison with Different Quantization Frameworks.
Method Data Time & Memory Low-bit
Free Efficient Compatible
PTQ
QAT
Ours
2.2 Diffusion Model Quantization
Model quantization is a dominant technique for optimizing the inference memory and speed of deep learning models by reducing the precision of the tensors used in computation. Few researches were done for quantizing diffusion models and they fall into two mainstream frameworks: Quantization-Aware Training (QAT) (Jacob et al., 2017; Li et al., 2022; Xu et al., 2023; Li et al., 2023b) and Post-Training Quantization (PTQ) (Wang et al., 2020; Nahshan et al., 2021; Li et al., 2021; Wei et al., 2023a; Liu et al., 2023). QAT methods such as Q-DM (Li et al., 2023b) consider value precision as an additional optimization constraint, training all parameters from scratch with the target task objective. While it is effective for low-bit quantization, it is even more resource-intensive than training a full-precision model. Efficient-DM (He et al., 2023a) further uses Low-rank adapters (LoRA) (Hu et al., 2021) to reduce training cost. However, it introduces extra weight parameters and still requires substantial training iterations.
Post-Training Quantization methods aim to calibrate the quantization parameters based on a small calibration set. Pioneering works such as PTQ4DM (Shang et al., 2023) and Q-Diffusion (Li et al., 2023a) focus on sampling the outputs of the full-precision model under different time steps with certain probabilities to construct the calibration set. ADP-DM (Wang et al., 2023) partition time steps into different groups and optimize each group with differentiable search algorithms. PTQ-D (He et al., 2023b) decomposes the quantization noise and correct them individually with full-precision model statistics. However, these methods generally fail under 4-bit or lower bit-width. To enable low-bit compatibility with high efficiency, our method adopts the general framework of PTQ but further introduces an efficient weight fine-tuning strategy. Table 1 shows the unique capabilities of our method, where we are able to realize QAT ability with PTQ efficiency.
@@ 3.1 Preliminaries
where $\hat{x}$ is the quantized integer result, $\text{round}(\cdot)$ represents rounding algorithms such as the round-to-nearest operator (Li et al., 2021) and AdaRound (Nagel et al., 2020), $s$ is referred to as the scaling factor and $Z$ is the zero-point. clamp is the function that clamps values into the range of $[q_{min},q_{max}]$, which is determined by the bit-width. Reversely, transforming the quantized values back into the full-precision form is formulated as:
@@
This is denoted as the dequantization process. The quantization and dequantization processes are performed on both model weights and layer outputs (also termed as ’activation’). Equation 2 indicates that the quantization error is composed of two factors: The clipping error produced by range clamping and the rounding error caused by the rounding function, where they exhibit a trade-off relationship (Li et al., 2021). While previous approaches strive for an optimal balance between the two errors, they neglect the intrinsic mechanisms in quantized diffusion models. In the next section, we investigate three properties of diffusion models that relate to quantization performance and show why typical quantization methods tend to fail.
@@ 3.2 Quantization-aware Properties of Diffusion Models
Property : The activation outputs are likely to have a majority of values close to zero, with the other values being numerically large and important. Previous works (Shang et al., 2023; Li et al., 2023a) paid attention to the varying activation distribution over different time steps, but none of them focused on the activation distribution itself. As shown in Figure 2, we find that in some layers, though the majority of values are close to 0, there exists values that are relatively large and diverse. Take the bin plot on the left for example, the original activation values (blue line) ranges from [-10, 34] but with over 99% values between [-0.6, 1.7], posing difficulties in minimizing the clipping error. This phenomenon is different from the outlier distribution observed in LLMs (Lin et al., 2023; Xiao et al., 2023), where the outliers are always positive and show up at certain positions consistently (Wei et al., 2023b). Moreover, we observe that the larger values are important for generation performance. With the large values being important and the small values appearing frequently, neither of them are negligible and need to be carefully quantized at the same time. Unfortunately, typical quantization methods fall short of this ability under low-bit settings, where the rounding error often outweighs the clipping error during optimization and results in over-clipped large values, generating corrupted images. This phenomenon inspires us to reshape the activation distributions via model finetuning to attain more quantization-friendly distributions.
@@
Figure 2: Illustrations of imbalanced activation distributions. We cluster activation values into uniformly distributed bins and manipulate the frequency for better visualization. In the activation distribution of the full-precision model, the majority of values cluster near zero with sporadic large values, presenting challenges for low-bit quantization. Our method reduces the large values to enable easier quantization for improved performance.
@@
Property : Accurate temporal information is critical for quantized diffusion model performance. As shown in Equation 1, diffusion model predictions are highly correlated with time step information. Specifically, time steps are transformed and projected into time embeddings and then added onto the image features, guiding the model to denoise correctly. Table 2 shows the performance comparison of quantizing time embeddings or preserving their full-precision values in the LDM-4 of LSUN-Bedrooms (Yu et al., 2015). Under W8A8 and W4A8, quantizing the time embeddings leads to an increase of 0.81 and 1.04 (relatively 15%) in FID, respectively. This indicates that inaccurate time embedding quantization can lead to large degradation in generation performance, yielding even larger corruptions under lower bit-widths. Due to the different focus diffusion models have on image generation at different time steps, we infer that inaccurate time embeddings can cause mismatched input image and model functionality, resulting in possible oscillations in the sequence of noise removal. Thus, preserving diffusion model performance under quantization requires accurate temporal information.
@@
Property : Different activations have different sensitivity towards decreasing bit-width. Low-bit activation quantization is often more difficult than weight quantization and remains the key challenge for diffusion model quantization. Compared with weights than can be categorized into linear layers and convolutional layers, activation types are more diverse and complicated. We hence examine the effects on model performance when quantizing different activations. Figure 3 shows how sensitive different activations are to quantization in Stable Diffusion. We quantize three different types of activations to lower bits while maintaining the others’ bit-width (8-bit). We observe that the FeedForward layer (Feng et al., 2023) activations cause generation failure at 6 bits, whereas the activations of all other linear layers (containing 5 times more layers) barely fail at 4 bits and all convolution layers (containing 3 times more layers) only fail at 4 bits. This phenomenon indicates that a small amount of activations are especially sensitive to low-bit quantization, making them essential to be specially dealt with.
Table 2: FP & Quantized Time Embedding Comparisons.
Settings FID $\downarrow$ sFID $\downarrow$
W8A8 (Quantized Time Embedding) 7.58 22.07
W8A8 (FP Time Embedding) 6.77 22.03
W8A8 (TAQuant+$\text{SLA}_{\text{TE}}$) 5.61 21.22
W4A8 (Quantized Time Embedding) 8.59 22.74
W4A8 (FP Time Embedding) 7.55 21.69
W4A8 (TAQuant+$\text{SLA}_{\text{TE}}$) 6.95 23.17

### v3

@@ Abstract
The practical deployment of diffusion models still suffers from the high memory and time overhead. While quantization paves a way for compression and acceleration, existing methods unfortunately fail when the models are quantized to low-bits. In this paper, we empirically unravel three properties in quantized diffusion models that compromise the efficacy of current methods: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules. To alleviate the intensified low-bit quantization difficulty stemming from the distribution imbalance, we propose finetuning the quantized model to better adapt to the activation distribution. Building on this idea, we identify two critical types of quantized layers: those holding vital temporal information and those sensitive to reduced bit-width, and finetune them to mitigate performance degradation with efficiency. We empirically verify that our approach modifies the activation distribution and provides meaningful temporal information, facilitating easier and more accurate quantization. Our method is evaluated over three high-resolution image generation tasks and achieves state-of-the-art performance under various bit-width settings, as well as being the first method to generate readable images on full 4-bit (i.e. W4A4) Stable Diffusion. Code is available here.
@@ 1 Introduction
Diffusion models [6, 15, 32] have achieved great success in the field of image generation recently. Given the high capacity and flexibility of diffusion models, there still remains two challenges that limit their efficiency [4]. The first obstacle is the time-consuming denoising process which requires hundreds to thousands of inference time steps that slows down the generation speed drastically. The other one is the increasing model size driven by the need of better image fidelity and higher image resolutions. Both factors contribute to considerable latency and increased computational requirements, impeding the application of diffusion models to real-world settings where both time and computational power are carefully restricted.
@@
Neural network quantization offers a feasible solution for accelerating inference speed and reducing memory consumption simultaneously [9]. It aims to compress high-bit model parameters into low-bit approximations with negligible performance degradation. For example, 4-bit weight and 4-bit activation quantization can achieve up to 8$\times$ inference time speedup and 8$\times$ memory reduction theoretically [23]. The benefits will increase consistently when the bit-widths are further lowered. Hence, low-bit quantization of diffusion models emerges as a viable approach for efficiency enhancement.
@@
However, existing diffusion model quantization approaches fail under low-bit settings. These methods focus either on timestep-aware calibration data construction [19, 34] or quantization noise correction [12], targeting at tailoring existing quantization techniques to the properties of diffusion models that are distinct from other model types such as CNNs [31] and ViTs [2]. Such approaches overlook the intrinsic diffusion model mechanisms that are related with quantization, leading to inconsistencies between method deployment and model characteristics. Therefore, we intend to identify unique features in diffusion models that are associated with quantization.
@@
In this paper, we reveal three properties of quantized diffusion models that hinder effective quantization: The activation distributions are likely to be imbalanced, where the majority of values are close to 0 but the important and large values are rare and appear inconsistently. Conventional quantization methods tend to either approximate the large and sparse values, inadequately estimating the numerous small values, or focus on small values while overlooking the large ones, thereby impeding the reduction of quantization error. Diffusion models exhibit varying functions at distinct time steps [3], such as forming object outlines at early time steps and constructing details latter. Thus, preserving accurate temporal information is important during quantization. However, current methods struggle to represent extensive time steps with limited bit-width, leading to similar time embeddings across different time steps and results in error accumulation. Diffusion models typically have complex network architectures, incorporating convolutional layers and linear layers that construct various kinds of attention blocks [7] and residual blocks [10]. Whereas previous works consider each model component as equally important and apply quantization uniformly, we reveal that certain layers are particularly sensitive to perturbations from quantization, while others are more resilient.
Table 1: Comparison with different frameworks.
Method Data Time & Memory Low-bit Fully Free Efficient Compatible Quantized PTQ [19] QAT [20] EfficientDM [11] Ours
@@
Addressing the revealed properties, we propose a novel quantization approach for diffusion models, termed QuEST (Quantization via Efficient Selective FineTuning). Confronting property , we first theoretically justify that weight finetuning can enhance model robustness towards large perturbations in activations under low-bit settings for reducing quantization error, whereas previous methods struggle to find appropriate quantization parameters. Then we empirically show that by finetuning the model, the activation distributions are actually modified to be more amenable to quantization by reducing the significant values (large yet sparse), as shown in Figure 1(a). Following the idea of weight finetuning, we compare different bit-width models (Figure 1(b)) and identify two types of layers that are responsible for performance degradation: Time embedding layers exhibiting property and attention-related layers corresponding to property . Thus, we selectively and progressively finetune these small layer subsets in conjunction with all activation quantization parameters, as illustrated in Figure 1(c). With less than 10% of the total parameters involved, QuEST not only substantially enhances low-bit quantized model performance, but is also notably time-efficient and requires only supervision from the full-precision model in a data-free manner. Our contributions are summarized as follows:
@@
  • We uncover and validate three properties that present challenges to low-bit diffusion model quantization: imbalanced activation distributions, imprecise temporal information, and vulnerability to perturbations of specific modules. We also theoretically discuss why previous methods are expected to fail and justify the sufficiency of weight finetuning under low-bit settings.
@@
  • Addressing the identified properties, we introduce QuEST, a new finetuning strategy that trains the model selectively and progressively, achieving low-bit quantization capability with high time and memory efficiency. An efficient activation quantizer is also proposed for time information estimation. Our method can be executed in a data-free and parameter-efficient manner.
@@
The detailed related works on diffusion model inference and quantization are referred to appendix for space restriction. Table 1 is a compact summary that shows the unique capabilities of our method, where we are able to overcome the limitations of previous quantization methods. In comparison with EfficientDM [11], we include a brief discussion here: Technically, EfficientDM aims to finetune via low-rank weight representations [16]. However, it introduces additional parameters for each layer and requires extensive iterations of training on LoRA weights. Our method instead does not include additional parameters and only needs to finetune the model for limited iterations with the same batch size. We have also found that EfficientDM does not quantize the matrix multiplications in the attention mechanism, while we quantize the matrices to 8-bits. Moreover, EfficientDM always keeps the first three layers and the last layer as full 8-bit (weight and activation), as well as keeping skip-connections and some linear layers as full-precision during implementation. Our method instead only kept the first and last layer as 8-bit and quantize all other layers to the required bit-width.
Figure 1: Overview of our findings and method. (a) Illustration of property and a way to mitigate the quantization difficulty. (b) Demonstration of how property and property are identified via comparing different bit-width models. (c) Framework of the proposed method at time step $t$. To address the quantization difficulty and achieve more quantization-friendly activation distribution efficiently, we reduce significance by selectively and progressively updating partial model layers’ weights. $e(t)$ and $x_{t}$ are the input time embedding and input image representation respectively. While $\mathbf{W}_{\text{TE}}$ and $\mathbf{W}_{\text{A}}$ are selected layer weights to be optimized using Equation 2.3.2, $\mathbf{W}_{l^{{}^{\prime}}}$ and $\mathbf{W}_{l^{{}^{\prime\prime}}}$ are frozen during finetuning. More details are included in Section 2.3.2.
@@ 2.1 Preliminaries
where $\hat{x}$ is the quantized integer result, $\text{round}(\cdot)$ represents rounding algorithms such as the round-to-nearest operator [22] and AdaRound [27], $s$ is referred to as the scaling factor and $Z$ is the zero-point. clamp is the function that clamps values into the range of $[q_{min},q_{max}]$, which is determined by the bit-width. Reversely, transforming the quantized values back into the full-precision form is:
@@
This is denoted as the dequantization process. The quantization and dequantization processes are performed on both model weights and layer outputs (also termed as ’activation’). Equation 1 indicates that the quantization error is composed of two factors: The clipping error produced by range clamping and the rounding error caused by the rounding function, where they exhibit a trade-off relationship [22]. While previous approaches strive for an optimal balance between the two errors, they neglect the intrinsic mechanisms in quantized diffusion models. In the next section, we investigate three properties of diffusion models that relate to quantization performance and show why typical quantization methods tend to fail.
@@ 2.2 Quantization-aware Properties of Diffusion Models
Property : The activation outputs are likely to have a majority of values close to zero, with the other values being numerically large and important. Previous works [19, 34] paid attention to the varying activation distribution over different time steps, but none of them focused on the activation distribution itself. In Figure 2, we cluster activation values into uniformly distributed bins and manipulate the frequency for better visualization. We find that in some layers, though the majority of values are close to 0, there exists values that are relatively large and diverse, which we term as "significant values" (circled in green). Take the bin plot on the left for example, the original activation values (blue line) ranges from [-10, 34] but with over 99% values between [-0.6, 1.7]. More visualizations can be found in Figure 6. This phenomenon poses difficulties in minimizing the clipping error and is unfriendly for effective quantization.
@@
Figure 2: Illustration of imbalanced activation distributions. In the activation distribution of the full-precision model, the majority of values cluster near zero with sporadic large values, presenting challenges for low-bit quantization. Our method reduces the significant values to enable easier quantization for improved performance.
Moreover, we observe that these significant values are important for generation performance preservation (see Figure 7). With the significant values being important and the small values appearing frequently, neither of them are negligible and need to be carefully quantized at the same time. Unfortunately, typical quantization methods fall short of this ability under low-bit settings, where the rounding error often outweighs the clipping error during optimization and results in over-clipped significant values, generating corrupted images. This inspires us to reshape the activation distributions via model finetuning to attain more quantization-friendly distributions, as depicted in Figure 1(a).
Table 2: Time embedding type comparison.
Bit-width Time Embedding FID $\downarrow$ sFID $\downarrow$ W8A8 PTQ Quantized 7.58 22.07 Full Precision 6.77 22.03 Ours 5.61 21.22 W4A8 PTQ Quantized 8.59 22.74 Full Precision 7.55 21.69 Ours 6.95 23.17
@@
Property : Accurate temporal information is critical for quantization performance. Diffusion model predictions are highly correlated with time step information. Specifically, time steps are transformed and projected into time embeddings and then added onto the image features, guiding the model to denoise correctly. Table 2 shows the performance comparison of quantizing time embeddings or preserving their full-precision values in the LDM-4 of LSUN-Bedrooms [39]. Under W8A8 and W4A8, quantizing the time embeddings leads to an increase of 0.81 and 1.04 (relatively 15%) in FID, respectively. This indicates that inaccurate time embedding quantization can lead to large degradation in generation performance, yielding even larger corruptions under lower bit-widths. Due to the different focus diffusion models have on image generation at different time steps, we infer that inaccurate time embeddings can cause mismatched input image and model functionality, resulting in possible oscillations in the sequence of noise removal. Thus, preserving diffusion model performance under quantization requires accurate temporal information.
Figure 3: Effect of decreasing different activations’ bit-width on the model performance. The generation failure of FeedForward layers emerges at 6 bits, while all other linear layers barely fail at 4 bits and all convolution layers only fail at 4 bits.
@@
Property : Different activations have different sensitivity towards decreasing bit-width. Low-bit activation quantization is often more difficult than weight quantization and remains the key challenge for diffusion model quantization. Compared with weights that can be categorized into linear layers and convolutional layers, activation types are more diverse and complicated. Figure 3 shows the sensitivity of different layers’ activations to quantization. Specifically, we quantize three different types of activations to lower bits while maintaining the others’ bit-width to 8-bit, and observe how the decreasing bit-width effects generation performance. Surprisingly, we observe that the FeedForward layer [8] activations cause generation failure at as large as 6 bits, whereas the activations of all other linear layers (containing 5 times more layers) barely fail at 4 bits and all convolution layers (containing 3 times more layers) only fail at 4 bits. This indicates that a small amount of activations are especially sensitive to low-bit quantization, making them essential to be specially dealt with.